A soft drink industry wants to make orange drink with a 40% content of orange juice. It has two types of orange drink: The first contains 80% pure juice and the second 20%. How many liter to mix from the second kind with 1000 liter of the first
Found 2 solutions by Theo, richwmiller:
Answer by Theo(13342)

1000 = number of liters of the first juice.
y = number of liters of the second juice.

total number of liters of both juices = 1000 + y.

first juice has 80% pure orange juice.
second juice has 20% pure orange juice.

you want to have a combined mixture of 40% pure orange juice.

.8 * 1000 + .2 * y = .4 * (1000 + y)

simplify to get 800 + .2 * y = 400 + .4 * y

subtract 400 from both sides of this equation and subtract .2 * y from both sides of this equation to get:

400 = .2 * y

divide both sides of this equation by .2 and solve for y to get:

y = 400 / .2 = 2000.

1000 liters of 80% pure orange juice contains 800 liters of pure orange juice.
2000 liters of 20% pure orange juice contains 400 liters of pure orange juice.

total liters of juice is 3000
total liters of contained pure orange juice is 1200 liters.
1200 / 3000 = .4 = 40%

solution is confirmed to be good.

solution is 2000 liters of the second juice needs to be mixed with 1000 liters of the first juice.

Answer by richwmiller(17219)

Orange juice is always 100% orange juice.
As soon as you dilute it is not orange juice.
Anything else is orange drink or punch.